OpenStudy (anonymous):

what is 4 consecutive even number that added up to 116

OpenStudy (anonymous):

even numbers have 0,2,4,6, or 8 for the last digit. so just do trial an error i guess.

OpenStudy (anonymous):

well first, we can represent that first consecutive even number with x. since we are dealing with even numbers, add by multiples of 2. x=first EVEN number x+2= second x+4= third x+6= fourth we take these and add them together and set it equal to 116 x+(x+2)+(x+4)+(x+6)=116 4x+12=116 x=26 26 is the first EVEN consecutive number. plug this value to get the rest.

OpenStudy (nisha_priya05):

take x, x+2,x+4,x+6 as 4 consecutive even nos. add them and put =116 find the value of x and substitute. you will get your ans.